# An Ising Similarity Regression Model

### v1

@@
Abstract: Understanding the dependence structure between response variables is an important component in the analysis of correlated multivariate data. This article focuses on modeling dependence structures in multivariate binary data, motivated by a study aiming to understand how patterns in different U.S. senators’ votes are determined by similarities (or lack thereof) in their attributes, e.g., political parties and social network profiles. To address such a research question, we propose a new Ising similarity regression model which regresses pairwise interaction coefficients in the Ising model against a set of similarity measures available/constructed from covariates. Model selection approaches are further developed through regularizing the pseudo-likelihood function with an adaptive lasso penalty to enable the selection of relevant similarity measures. We establish estimation and selection consistency of the proposed estimator under a general setting where the number of similarity measures and responses tend to infinity. Simulation study demonstrates the strong finite sample performance of the proposed estimator in terms of parameter estimation and similarity selection. Applying the Ising similarity regression model to a dataset of roll call voting records of 100 U.S. senators, we are able to quantify how similarities in senators’ parties, businessman occupations and social network profiles drive their voting associations. Key words and phrases: Conditional dependence, Ising model, Lasso, Model selection, Multivariate data, Pseudo-likelihood
@@ 1 Introduction
The study of the dependence structure in correlated multivariate data has drawn much attention in recent years, due to the increasing accessibility of multi-response data across many disciplines such as finance (Tsay, 2013), economics (McElroy and Trimbur, 2023) and ecology (Hui, Müller, and Welsh, 2023). For a multivariate response vector $\bm{y}=(y_{1},\cdots,y_{p})^{\top}$, the dependencies across any set of two components $y_{j}$ and $y_{j^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ can be modeled by specifying a joint probabilistic distribution for $\bm{y}$. Two prominent examples are Gaussian graphical models (Whittaker, 1990) and Ising models (Ising, 1925), which specify a probability density function (pdf) and a probability mass function (pmf) for a continuous and binary response vector $\bm{y}$, respectively, where the pdf and pmf of these two models both contain a component of the form $\sum_{1\leq j<j^{\prime}\leq p}\exp(\theta_{jj^{\prime}}y_{j}y_{j^{\prime}})$ with $\theta_{jj^{\prime}}$ denoting a coefficient for the interaction term between $y_{j}$ and $y_{j^{\prime}}$. Accordingly, the set of $\theta_{jj^{\prime}};j,j^{\prime}=1,\cdots,p$} are referred to as pairwise interaction coefficients in these joint models. Moreover, in both models each $\theta_{jj^{\prime}}$ is directly tied to the conditional dependence relationship between $y_{j}$ and $y_{j^{\prime}}$ given the remaining elements in the response vector $\bm{y}$ (Hammersley-Clifford equivalence, Hammersley and Clifford, 1971): if $\theta_{jj^{\prime}}$ is close to/far away from zero, the conditional dependence between $y_{j}$ and $y_{j^{\prime}}$ is weak/strong, with the limiting case of $\theta_{jj^{\prime}}=0$ implying conditional independence. This interpretation of the pairwise interaction coefficient $\theta_{jj^{\prime}}$ has contributed to the popularity of Gaussian graphical models and Ising models for correlated multivariate data analysis; see for instance Meinshausen and Bühlmann (2006), Yuan and Lin (2007) and Friedman, Hastie, and Tibshirani (2007) for the Gaussian graphical models, and Majewski, Li, and Ott (2001), Guo et al. (2015) and Bhattacharya and Mukherjee (2018) for the Ising models, among many others. In this article, we focus on modeling the dependence relationships for a multivariate binary response vector $\bm{y}$ through, or equivalently, modeling the pairwise interaction coefficients $\theta_{jj^{\prime}}$ in, the Ising model. As demonstrated above, the conditional dependence of multivariate binary responses in the Ising model is captured by the pairwise interaction coefficients $\theta_{jj^{\prime}}$, which can be collected into a symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ where $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. A variety of approaches have been proposed to estimate and regularize $\bm{\Theta}$. For example, Banerjee, El Ghaoui, and d’ Aspremont (2008) proposed a block-coordinate descent algorithm to solve an approximate sparse maximum likelihood problem for the estimation of $\bm{\Theta}$. Ravikumar, Wainwright, and Lafferty (2010) considered a neighborhood estimation method based on fitting a separate regularized logistic regression with the lasso penalty to each binary response, while Höfling and Tibshirani (2009) and Guo et al. (2010) used a pseudo-likelihood function coupled with a lasso-type penalty to simultaneously estimate and regularize all pairwise interaction coefficients in $\bm{\Theta}$. Further penalized likelihood approaches for sparse $\bm{\Theta}$ estimation have been developed by Lee, Ganapathi, and Koller (2006) and Xue, Zou, and Cai (2012), among others. In contrast to estimating and regularizing the elements of the interaction matrix $\bm{\Theta}$, this article proposes a novel Ising similarity regression model which regresses the pairwise interaction coefficients $\theta_{jj^{\prime}}$ against a set of pairwise similarity measures $w_{jj^{\prime}}^{(k)}$; that is,
@@
where $\alpha_{k}$ denotes the regression coefficient associated with the $k$-th similarity measure $w_{jj^{\prime}}^{(k)}$ for $k=1,\cdots,K$. It is worth noting $w_{jj^{\prime}}^{(k)}$ measures the similarity between $j$ and $j^{\prime}$, which can be either observed directly as part of the data collection process, or induced from available auxiliary information. As a motivating example, we consider U.S. Senate roll call voting data where the binary response $y_{j}$ represents the $j$-th senator’s voting record (Yea or Nay) to a particular piece of legislation or bill. The pairwise interaction coefficient $\theta_{jj^{\prime}}$ thus captures the conditional dependence relationship between the $j$-th and $j^{\prime}$-th senators’ voting patterns, while $w_{jj^{\prime}}^{(k)}$ are constructed based on additional attributes of the $j$-th and $j^{\prime}$-th senators, such as political parties, ages, occupations and social network profiles. In such an example, the regression coefficients $\alpha_{k}$ in model (1.1) offer a clear, explicit quantification of how the $k$-th similarity measure between the $j$-th and $j^{\prime}$-th senators affects their conditional dependence when it comes to voting patterns.
@@
At this point, it is important to acknowledge the work of Cheng et al. (2014), who proposed an Ising regression model to regress the pairwise interaction coefficients $\theta_{jj^{\prime}}$ on a covariate vector $\bm{x}$. Such a model with $\theta_{jj^{\prime}}=\theta_{jj^{\prime}}(\bm{x})$ does not provide the same interpretation for our motivating example however, as their interaction coefficients $\theta_{jj^{\prime}}$ for different pairs $(j,j^{\prime})$ only depend on the same covariate vector $\bm{x}$ instead of the pairwise similarity measures $w_{jj^{\prime}}^{(k)}$ considered in model (1.1).
@@
Apart from the aforementioned example in political science, the proposed Ising similarity regression model has a multitude of applications across finance and ecology. For example, in the study of capital markets, $y_{j}$ can be an indicator denoting whether the $j$-th firm has distributed dividends to its shareholders, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th firms’ financial fundamentals such as market value, cash flow, and leverage. Also, in ecology, $y_{j}$ can be a binary variable indicating the presence or absence of the $j$-th species, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th species’ traits including but not limited to breeding season and overwintering behavior. For both of these examples, $\alpha_{k}$ in model (1.1) quantifies the effect of the $k$-th similarity measure on the conditional dependence relationships between the responses of the $j$-th and ${j^{\prime}}$-th companies or species. More broadly, the use of similarity measures in regression is also motivated by recent developments in covariance regression modeling (Zou et al., 2017, 2020, 2022), where the covariance between continuous responses is modeled as a linear combination of similarity measures. To the best of our knowledge though, the proposed Ising similarity regression model is the first to explicitly link similarity measures to the pairwise interaction coefficients (conditional dependence) of binary responses.
@@
To estimate the Ising similarity regression model, we propose a regularized pseudo-likelihood approach which augments the pseudo-likelihood function of model (1.1) with an adaptive lasso penalty (Zou, 2006). Doing so induces sparse estimation of the regression coefficients $\{\alpha_{k};k=1,\cdots,K\}$, which allows us to recover similarity measures that are truly relevant in explaining the conditional dependence relationships between the binary responses. Under a setting where the number of regression coefficients $K$ and responses $p$ tend to infinity, we establish estimation and selection consistency for the regularized pseudo-likelihood estimator. To select the tuning parameter in the adaptive lasso penalty, we employ a cross-validation approach which preserves the dependencies between the elements of $\bm{y}$. Simulation results support the theoretical findings of the proposed estimator, demonstrating its strong finite sample estimation and model selection performance. We apply the Ising similarity regression model to roll call voting records of 100 U.S. senators from the 117-th Congress, with results demonstrating how similarities of senators’ attributes and social network activities drive the association between their voting patterns. In particular, aside from the expected findings such as senators from the same state or party being more likely to vote similarly, we find that senators who are businessmen or share certain follower-followee relationships on Twitter tend to exhibit more similar voting patterns.
@@
The rest of this article is organized as follows. Section 2 introduces the Ising similarity regression model along with the proposed regularized pseudo-likelihood estimator. Section 3 discusses the theoretical properties of the regularized estimator. Section 4 presents simulation studies, while an application to U.S. roll call voting dataset is provided in Section 5. Section 6 offers some concluding remarks. All theoretical proofs of the theorems developed in this article are presented in the supplementary material.
@@ 2.1 Model Set-Up
where $\bm{\theta}=(\theta_{11},\cdots,\theta_{1p},\cdots,\theta_{p-1,p-1},\theta_{p-1,p},\theta_{pp})^{\top}$ is a $p(p+1)/2$-dimensional parameter vector, and the partition function $Z(\bm{\theta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp(\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}\theta_{jj^{\prime}}u_{j}u_{j^{\prime}})$ is an intractable normalization constant in the pmf. The main effect parameters in (2.2) are given by $\theta_{jj}$ for $j=1,\cdots,p$, while the pairwise interaction coefficients in the pmf are given by $\theta_{jj^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ and $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. As discussed in Section 1, each of the pairwise interaction coefficients $\theta_{jj^{\prime}}$ is related to the conditional dependence between $y_{j}$ and $y_{j^{\prime}}$, conditional on the other elements of the response vector $\bm{y}$ (Hammersley and Clifford, 1971). This makes the Ising model useful in studying the conditional dependence structure among binary responses, through analyzing the symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$.
Suppose now that, in addition to observing the binary response vector $\bm{y}$, we also record a set of $p\times p$ symmetric similarity matrices $\bm{W}_{k}=(w_{jj^{\prime}}^{(k)})_{p\times p}$ for $k=1,\cdots,K$, which may be available directly as part of the data collection process or constructed from auxiliary information variables $z_{j1},\cdots,z_{jK}$ associated to the $j$-th response for $j=1,\cdots,p$. For instance, in the motivating U.S. senators’ voting records data, one of the similarity matrices $\bm{W}_{k}$ is set as the $p\times p$ network adjacency matrix based on the follower-followee relationships on Twitter, while other similarity matrices are constructed based on auxiliary information variables $z_{jk}$ consisting of the $j$-th senator’s demographic attributes, such as political party, age, and occupation (see Section 5 for details). In the latter, each element $w_{jj^{\prime}}^{(k)}$ of the similarity matrix $\bm{W}_{k}$ measures the similarity between $z_{jk}$ and $z_{j^{\prime}k}$ for $j\neq j^{\prime}$ e.g., if $z_{jk}$ is quantitative then we set $w_{jj^{\prime}}^{(k)}=\exp(-|z_{jk}-z_{j^{\prime}k}|^{2})$, whereas if $z_{jk}$ is qualitative then $w_{jj^{\prime}}^{(k)}=1$ if $z_{jk}$ and $z_{j^{\prime}k}$ have the same categorical level (e.g., same occupation), otherwise, $w_{jj^{\prime}}^{(k)}=0$ (see also Johnson and Wichern, 1992). For completeness and reasons of parameter identifiability, the diagonals $w_{jj}^{(k)}$ are set to be zeros for all $j=1,\cdots,p$ and $k=1,\cdots,K$.
Given a set of similarity matrices $\bm{W}_{k}$ for $k=1,\cdots,K$, the Ising similarity regression model as introduced in equation (1.1) can be equivalently formulated as modeling the interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ via the form
@@
In equation (2.3), the vector of regression coefficients $\bm{\alpha}$ describes how similarities directly affect the pairwise interaction coefficients (conditional dependence relationships) of the binary responses. Note also that by utilizing similarity matrices to model the interaction matrix $\bm{\Theta}$, model (2.4) involves a substantially smaller number of parameters ($p+K$ parameters) compared to the standard Ising model in (2.2), which has $(p+1)p/2$ parameters. The proposed Ising similarity regression model is thus particularly useful when the dimension of the binary response vector $p$ is large: even when the number of similarity matrices $K$ grows at the same rate as the dimension $p$, the number of parameters in model (2.4) only grows linearly in $p$ compared to $O(p^{2})$ parameters in model (2.2).
Equation (2.3) is related to a variety of methods developed for linking a matrix parameter to a linear combination of known symmetric matrices; see, for example, Anderson (1973) and the subsequent research by Zwiernik, Uhler, and Richards (2017), and Zou et al. (2017, 2020, 2022) among others. Modeling the interaction matrix $\bm{\Theta}$ as a linear combination of the similarity matrices $\bm{W}_{k}$ for $k=1,\cdots,K$ and $\bm{\Delta}_{jj}$ for $j=1,\cdots,p$ in equation (2.3) is similar to linking a function of the covariance matrix to a linear combination of matrices (e.g., Chiu, Leonard, and Tsui, 1996; Pourahmadi, 1999; Bonat and Jørgensen, 2016), although to our knowledge this article is the first to establish such an idea for the Ising model. It is also worth noting that equation (2.3) includes the standard Ising model as a special case, by letting $\bm{W}_{1}=\bm{\Delta}_{12}+\bm{\Delta}_{21},\cdots,\bm{W}_{p-1}=\bm{\Delta}_{1p}+\bm{\Delta}_{p1},\bm{W}_{p}=\bm{\Delta}_{23}+\bm{\Delta}_{32},\cdots,\bm{W}_{p(p-1)/2}=\bm{\Delta}_{(p-1)p}+\bm{\Delta}_{p(p-1)}$, $\alpha_{1}=\theta_{12},\cdots,\alpha_{p-1}=\theta_{1p},\alpha_{p}=\theta_{23},\cdots,\alpha_{K}=\theta_{(p-1)p}$, and $K=p(p-1)/2$. As such, the Ising similarity regression model (2.4) not only allows us to estimate the interaction matrix by incorporating the similarities and/or auxiliary information, but also may be interpreted as a generalized presentation of the standard Ising model.
2.2 Estimation and Selection

### v2

# Supplementary

@@
Abstract: Understanding the dependence structure between response variables is an important component in the analysis of correlated multivariate data. This article focuses on modeling dependence structures in multivariate binary data, motivated by a study aiming to understand how patterns in different U.S. senators’ votes are determined by similarities (or lack thereof) in their attributes, e.g., political parties and social network profiles. To address such a research question, we propose a new Ising similarity regression model which regresses pairwise interaction coefficients in the Ising model against a set of similarity measures available/constructed from covariates. Model selection approaches are further developed through regularizing the pseudo-likelihood function with an adaptive lasso penalty to enable the selection of relevant similarity measures. We establish estimation and selection consistency of the proposed estimator under a general setting where the number of similarity measures and responses tend to infinity. Simulation study demonstrates the strong finite sample performance of the proposed estimator, particularly compared with several existing Ising model estimators in estimating the matrix of pairwise interaction coefficients. Applying the Ising similarity regression model to a dataset of roll call voting records of 100 U.S. senators, we are able to quantify how similarities in senators’ parties, businessman occupations and social network profiles drive their voting associations.
Key words and phrases: Conditional dependence, Ising model, Lasso, Model selection, Multivariate data, Pseudo-likelihood
@@ 1 Introduction
The study of the dependence structure in correlated multivariate data has drawn much attention in recent years, due to the increasing accessibility of multi-response data across many disciplines such as finance (Tsay, 2013), economics (McElroy and Trimbur, 2023) and ecology (Hui, Müller, and Welsh, 2023). For a multivariate response vector $\bm{y}=(y_{1},\cdots,y_{p})^{\top}$, the dependencies across any set of two components $y_{j}$ and $y_{j^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ can be modeled by specifying a joint probabilistic distribution for $\bm{y}$. Two prominent examples are Gaussian graphical models (Whittaker, 1990) and Ising models (Ising, 1925), which specify a probability density function (pdf) and a probability mass function (pmf) for a continuous and binary response vector $\bm{y}$, respectively, where the pdf and pmf of these two models both contain a component of the form $\sum_{1\leq j<j^{\prime}\leq p}\exp(\theta_{jj^{\prime}}y_{j}y_{j^{\prime}})$ with $\theta_{jj^{\prime}}$ denoting a coefficient for the interaction term between $y_{j}$ and $y_{j^{\prime}}$. Accordingly, the set of $\theta_{jj^{\prime}}:j,j^{\prime}=1,\cdots,p$} are referred to as pairwise interaction coefficients in these joint models. Moreover, in both models each $\theta_{jj^{\prime}}$ is directly tied to the conditional dependence relationship between $y_{j}$ and $y_{j^{\prime}}$ given the remaining elements in the response vector $\bm{y}$ (Hammersley-Clifford equivalence, Hammersley and Clifford, 1971): if $\theta_{jj^{\prime}}$ is close to/far away from zero, the conditional dependence between $y_{j}$ and $y_{j^{\prime}}$ is weak/strong, with the limiting case of $\theta_{jj^{\prime}}=0$ implying conditional independence. This interpretation of the pairwise interaction coefficient $\theta_{jj^{\prime}}$ has contributed to the popularity of Gaussian graphical models and Ising models for correlated multivariate data analysis; see for instance Meinshausen and Bühlmann (2006), Yuan and Lin (2007) and Friedman, Hastie, and Tibshirani (2007) for the Gaussian graphical models, and Majewski, Li, and Ott (2001), Guo et al. (2015) and Bhattacharya and Mukherjee (2018) for the Ising models, among many others.
In this article, we focus on modeling the dependence relationships for a multivariate binary response vector $\bm{y}$ through, or equivalently, modeling the pairwise interaction coefficients $\theta_{jj^{\prime}}$ in, the Ising model. As demonstrated above, the conditional dependence of multivariate binary responses in the Ising model is captured by the pairwise interaction coefficients $\theta_{jj^{\prime}}$, which can be collected into a symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ where $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. A variety of approaches have been proposed to estimate and regularize $\bm{\Theta}$. For example, Banerjee, El Ghaoui, and d’ Aspremont (2008) proposed a block-coordinate descent algorithm to solve an approximate sparse maximum likelihood problem for the estimation of $\bm{\Theta}$. Ravikumar, Wainwright, and Lafferty (2010) considered a neighborhood estimation method based on fitting a separate regularized logistic regression with the lasso penalty to each binary response, while Höfling and Tibshirani (2009) and Guo et al. (2010) used a pseudo-likelihood function coupled with a lasso-type penalty to simultaneously estimate and regularize all pairwise interaction coefficients in $\bm{\Theta}$. Further penalized likelihood approaches for sparse $\bm{\Theta}$ estimation have been developed by Lee, Ganapathi, and Koller (2006) and Xue, Zou, and Cai (2012), among others.
In contrast to estimating and regularizing the elements of the interaction matrix $\bm{\Theta}$, this article proposes a novel Ising similarity regression model which regresses the pairwise interaction coefficients $\theta_{jj^{\prime}}$ against a set of pairwise similarity measures $w_{jj^{\prime}}^{(k)}$; that is,
@@
where $\alpha_{k}$ denotes the regression coefficient associated with the $k$-th similarity measure $w_{jj^{\prime}}^{(k)}$ for $k=1,\cdots,K$. It is worth noting $w_{jj^{\prime}}^{(k)}$ measures the similarity between $j$ and $j^{\prime}$, which can be either observed directly as part of the data collection process, or induced from available auxiliary information. As a motivating example, we consider U.S. Senate roll call voting data where the binary response $y_{j}$ represents the $j$-th senator’s voting record (Yea or Nay) to a particular piece of legislation or bill, and $w_{jj^{\prime}}^{(k)}$ are constructed based on additional attributes of the $j$-th and $j^{\prime}$-th senators, such as political parties and occupations. In such an example, the regression coefficients $\alpha_{k}$ in model (1.1) offer a clear, explicit quantification of how the $k$-th similarity measure between the $j$-th and $j^{\prime}$-th senators affects their conditional dependence when it comes to voting patterns.
@@
At this point, it is important to acknowledge the work of Cheng et al. (2014), who proposed an Ising regression model to regress the pairwise interaction coefficients $\theta_{jj^{\prime}}$ on a covariate vector $\bm{x}$. Such a model with $\theta_{jj^{\prime}}=\theta_{jj^{\prime}}(\bm{x})$ does not provide the same interpretation for our motivating example however, as their interaction coefficients $\theta_{jj^{\prime}}$ for different pairs $(j,j^{\prime})$ only depend on the same covariate vector $\bm{x}$ instead of the pairwise similarity measures $w_{jj^{\prime}}^{(k)}$ considered in model (1.1). Also, in a parallel line of research, several studies have considered using covariates to model the precision matrix encoding the dependence structure in Gaussian graphical models. Liu et al. (2010) partitioned the covariate space based on classification and regression trees into multiple subspaces that can have different precision matrices, while Lee and Xue (2018) proposed a nonparametric mixture of Gaussian graphical models with mixture probabilities and precision matrices that are both covariate-dependent. More recently, Wang et al. (2022) and Ni, Stingo, and Baladandayuthapani (2022) proposed Bayesian approaches to model the elements of the precision matrix as linear functions of covariates. As these studies focus on precision matrices of Gaussian graphical models, they are not directly applicable to the multivariate binary response setting which we focus on in this article.
@@
Apart from the aforementioned example in political science, the proposed Ising similarity regression model has a multitude of applications across finance and ecology. For example, in the study of capital markets, $y_{j}$ can be an indicator denoting whether the $j$-th firm has distributed dividends to its shareholders, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th firms’ financial fundamentals such as market value, cash flow, and leverage. Also, in ecology, $y_{j}$ can be a binary variable indicating the presence or absence of the $j$-th species, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th species’ trait values; see an application of the proposed model to such an ecology dataset in Section LABEL:sec:beetle_application of the supplementary material, to illustrate the wide applicability of the proposed method. More broadly, the use of similarity measures in regression is also motivated by recent developments in covariance regression modeling (Zou et al., 2017, 2020, 2022), where the covariance between continuous responses is modeled as a linear combination of similarity measures. Indeed, the proposed model can be written in a matrix regression form, and this connects to the wider literature linking (functions of) matrix parameters to a linear combination of matrices; see for example Anderson (1973), Pourahmadi (1999) and Bonat and Jørgensen (2016). To the best of our knowledge though, the proposed Ising similarity regression model is the first to establish such an idea for the Ising model specifically, by explicitly linking similarity matrices $\bm{W}_{k}=(w_{jj^{\prime}}^{(k)})_{p\times p}$ to the interaction matrix $\bm{\Theta}$ that captures the conditional dependence of binary responses.
To estimate the proposed model, we study a regularized pseudo-likelihood approach which augments the pseudo-likelihood function of model (1.1) with an adaptive lasso penalty (Zou, 2006). Doing so induces sparse estimation of the regression coefficients $\{\alpha_{k}:k=1,\cdots,K\}$, which allows us to recover similarity measures that are truly relevant in explaining the conditional dependence relationships between the binary responses. It is important here to highlight that this article differs from the aforementioned studies on Ising model estimation, as our main focus is to induce sparsity on the regression coefficients $\alpha_{k}$ as opposed to the similarity measures $w_{jj^{\prime}}^{(k)}$ and the resulting interaction matrix $\bm{\Theta}$. Put another way, we aim to identify important drivers of the conditional dependence relationships i.e., similarity selection rather than edge selection, by treating the similarity measures as given covariates.
@@
Under a setting where the number of regression coefficients $K$ and responses $p$ tend to infinity, we establish estimation and selection consistency for the regularized pseudo-likelihood estimator. To select the tuning parameter in the adaptive lasso penalty, we employ a cross-validation approach which preserves the dependencies between the elements of $\bm{y}$. Simulation results support the theoretical findings of the proposed estimator, demonstrating its strong finite sample estimation and model selection performance. Specifically, the proposed estimator not only outperforms other estimators such as the unpenalized estimator and the lasso-penalized estimator in estimating the parameters of the Ising similarity regression model, but also performs much better than the traditional Ising model estimators that ignore the additional information from similarity measures in estimating the Ising model interaction matrix $\bm{\Theta}$. Additionally, we carry out simulation studies to compare the similarity selection performance of our cross-validation approach to the use of AIC (Akaike, 1998) and BIC (Schwarz, 1978) criteria for choosing the tuning parameter, with results showing that BIC has a comparable performance to cross-validation approach while AIC tends to suffer from overfitting. We apply the Ising similarity regression model to roll call voting records of 100 U.S. senators from the 117-th Congress, with results demonstrating how similarities of senators’ attributes and social network activities drive the association between their voting patterns. In particular, aside from the expected findings such as senators from the same state or party being more likely to vote similarly, we find that senators who are businessmen or share certain follower-followee relationships on Twitter tend to exhibit more similar voting patterns.
@@
The rest of this article is organized as follows. Section 2 introduces the Ising similarity regression model along with the proposed regularized pseudo-likelihood estimator. Section 3 discusses the theoretical properties of the regularized estimator. Section 4 presents simulation studies, while an application to U.S. roll call voting dataset is provided in Section 5. Section 6 offers some concluding remarks. All theoretical proofs of the theorems developed in this article, along with detailed empirical comparisons to other estimation approaches, as well as an additional application to Scotland Carabidae ground beetle dataset, are presented in the supplementary material.
@@ 2.1 Model Set-Up
where $\bm{\theta}=(\theta_{11},\cdots,\theta_{1p},\cdots,\theta_{p-1,p-1},\theta_{p-1,p},\theta_{pp})^{\top}$ is a $p(p+1)/2$-dimensional parameter vector, and the partition function $Z(\bm{\theta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp(\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}\theta_{jj^{\prime}}u_{j}u_{j^{\prime}})$ is an intractable normalization constant in the pmf. The main effect parameters in (2.2) are given by $\theta_{jj}$ for $j=1,\cdots,p$, while the pairwise interaction coefficients in the pmf are given by $\theta_{jj^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ and $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. As discussed in Section 1, the symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ is useful for studying the conditional dependence structure among binary responses.
Suppose now that, in addition to observing the binary response vector $\bm{y}$, we also record a set of $p\times p$ symmetric similarity matrices $\bm{W}_{k}=(w_{jj^{\prime}}^{(k)})_{p\times p}$ for $k=1,\cdots,K$, which may be available directly as part of the data collection process or constructed from auxiliary information variables $z_{j1},\cdots,z_{jK}$ associated to the $j$-th response for $j=1,\cdots,p$. In the latter, each element $w_{jj^{\prime}}^{(k)}$ of the similarity matrix $\bm{W}_{k}$ measures the similarity between $z_{jk}$ and $z_{j^{\prime}k}$ for $j\neq j^{\prime}$. For instance, if $z_{jk}$ is quantitative, then we can set $w_{jj^{\prime}}^{(k)}=\exp(-|z_{jk}-z_{j^{\prime}k}|^{2})$, whereas if $z_{jk}$ is qualitative then we set $w_{jj^{\prime}}^{(k)}=1$ if $z_{jk}$ and $z_{j^{\prime}k}$ have the same categorical level, and $w_{jj^{\prime}}^{(k)}=0$ otherwise (see also Johnson and Wichern, 1992). For completeness and reasons of parameter identifiability, the diagonals $w_{jj}^{(k)}$ are set to be zeros for all $j=1,\cdots,p$ and $k=1,\cdots,K$.
Given a set of similarity matrices $\bm{W}_{k}$ for $k=1,\cdots,K$, the Ising similarity regression model as introduced in equation (1.1) can be equivalently formulated as modeling the interaction matrix $\bm{\Theta}$ via the form
@@
In equation (2.3), the vector of regression coefficients $\bm{\alpha}$ describes how similarities directly affect the pairwise interaction coefficients (conditional dependence relationships) of the binary responses. Note also by utilizing similarity matrices to model the interaction matrix $\bm{\Theta}$, model (2.4) involves a substantially smaller number of parameters ($p+K$ parameters) compared to the standard Ising model in (2.2), which has $(p+1)p/2$ parameters. The proposed Ising similarity regression model is thus particularly useful when the dimension of the binary response vector $p$ is large: even when the number of similarity matrices $K$ grows at the same rate as the dimension $p$, the number of parameters in model (2.4) only grows linearly in $p$ compared to $O(p^{2})$ parameters in model (2.2).
2.2 Estimation and Similarity Selection